# Cross Domain Early Crop Mapping using CropGAN and CNN Classifier

### v1

@@ III-B Reference Data
The CDL and the Canada AAFC Annual Crop Inventory are used as the ground truth for the source domains. The CDL [7], a crop-specific land cover raster map dataset available for the entire conterminous U.S. land area (CONUS) at 30 m resolution provided by the USDA, regularly provides information on the annual temporal and spatial distribution of corn, as well as the area dedicated to its cultivation. However, for the target domains, official ground reference data for China is currently unavailable. Fortunately, [6] published maps of corn and soybean with a spatial resolution of 10 meters for Northeast China from 2017 to 2019. As a result, the 2019 crop maps from [6] are regarded as the ground truth reference for the study area in China. Additionally, the Canada AAFC Annual Crop Inventory is used as the ground truth for Canada with a spatial resolution of 30 meters.
@@ III-C Remote Sensing Data
As described, our primary objective is to locate the specific crop (corn) with cross-domain problems at an earlier growth stage. As a result, the time series remote sensing data should start after the planting period and conclude before the onset of the crop harvest period. Referring to Figure 3, it can be observed that the corn harvesting season typically starts in early September in the USA and begins in early October in China and Canada. Planting, on the other hand, starts at the beginning of April in the USA, in mid-April in Jilin Province, China, and at the beginning of May in Québec, Canada. Furthermore, to minimize the impact of variations in the crop calendar, the beginning dates of our annual collection of remote sensing data for different countries are based on the average planting dates shown in Figure 3. For Jackson County the date spans from May 1st to July 30th, encompassing a total of nine time points with a ten-day observation window. The start date of the extracted remote sensing data from China is ten days earlier than that from the study areas in the USA. The extracted data for China encompasses MSI images from May 1st, while for Cavalier County, USA, it starts from May 10th. Similarly, for Canada, the data involves MSI images beginning on May 30th, a 30-day later from the study area in Sussex County, USA, which started on May 1st. This approach gets the remote sensing data which covers the vegetation growth session before the crop harvesting season.
@@ IV Methodology
IV-A System Overview
Figure 6: The architecture of our cross-domain crop mapping system.
The system architecture of our cross-domain crop mapping method is illustrated in Figure 6. The system comprises three key components: the pre-processor, the domain mapper, and the crop mapper. The pre-processor provides the composited MSI images and employs a linear interpolation method to fill gaps in the images, ensuring a complete time series of remote sensing data. The domain mapper, based on the adversarial neural network, facilitates the transformation of time-series remote sensing data from the target domain to the source domain. Lastly, the crop mapper, utilizing a CNN-based model, is trained on the labelled data from the source domain. Then it receives the transformed target domain data from the domain mapper as input data to map the crop of the interest in the target domain. This comprehensive system design enables robust and precise early crop mapping within the target domain, in scenarios where labelled data is scarce or absent, addressing the challenge of domain shift.
IV-B Pre-processor
@@
The pre-processor aims at providing complete time-series remote sensing data by filling gaps between MSI images due to cloud cover, atmospheric interference, or sensor limitations. The processing pipeline is outlined in Figure 7.
Initially, the pre-processor retrieves MSI images encompassing the entire specified study area. These images, sourced from the Sentinel-2 Dataset, are captured at regular intervals of 10 days as an image group. Ultimately, nine image groups are assembled, starting from the crop planting date.
Concurrently, to exclude non-agricultural lands in the MSI images, the pre-processor extracts the cropland mask during the crop growing season from the Dynamic World dataset and reprojects it into a consistent 30-meter resolution to ensure uniformity. After filtering the non-agricultural lands out using the cropland mask, the MSI images in nine image groups are specific to the agricultural lands.
Within each image group, the pre-processor employs Sentinel Hub’s cloud detector [30] to apply cloud masks on the Sentinel-2 MSI images. This process amalgamates the images, preserving the mean values of the cloud-free MSI images. Consequently, a composite remote sensing image at a clear, 30-meter resolution is generated.
Through the pre-processor, a set of nine high-quality time-series remote sensing images is generated. These images are characterized by their absence of clouds and gaps, and they consist of six bands, all with a spatial resolution of 30 meters. These images provide a comprehensive and continuous record of the vegetation dynamics over time, enabling an in-depth analysis of temporal patterns and changes in vegetation growth.
IV-C Domain Mapper
Figure 8: The CropGAN structure.
The domain mapper is referred as CropGAN, which stands for Crop Generative Adversarial Network. CropGAN serves as the key component in our framework for transforming time-series remote sensing data points between different domains. As shown in Figure 8, the domain mapper consists of two generator networks, $G$ and $F$, and two discriminator networks, $D_{X}$ and $D_{Y}$. The visual representations of the generator and discriminator structures can be observed in Table II and Table III, respectively. Given two crop mapping domains $X$ and $Y$, $\left\{x_{i}\right\}^{N}_{i=1}$ is the time-series remote sensing data in Domain $X$, and $\left\{y_{j}\right\}^{M}_{j=1}$ is the time-series remote sensing data in Domain $Y$. CropGAN is trained to learn the mappings between different data distributions from different domains: $G:X\rightarrow Y$ and $F:Y\rightarrow X$. In the training process, the generators and discriminators play distinct roles. The generator’s primary task is to transform the data from one domain to another, while the discriminator’s main objective is to differentiate between real and transformed data. As the training progresses, both networks improve iteratively until an equilibrium is reached where the generator generates highly realistic data and the discriminator cannot reliably distinguish between real and fake samples. For example, when the discriminator $D_{Y}$ can not distinguish whether the data generated by the generator $G$ is from source domain $Y$ or not, it means that the generator $G$ has been well-trained.
@@ IV-C Domain Mapper
Input 9x6x1 9x6x1 -
Encoder 1 9x6x1 7x5x4 LeakyReLU
Encoder 2 7x5x4 5x4x8 LeakyReLU
Encoder 3 5x4x8 3x3x16 LeakyReLU
Encoder 4 3x3x16 1x2x32 LeakyReLU
Decoder 4 1x2x32 3x3x16 LeakyReLU
Decoder 3 3x3x16 5x4x8 LeakyReLU
Decoder 2 5x4x8 7x5x4 LeakyReLU
Decoder 1 7x5x4 9x6x2 LeakyReLU
Decoder 1 7x5x2 9x6x1 LeakyReLU
Output 9x2x1 9x2x1 ReLU
TABLE III: The Discriminator Structure. IN represents an instance normalization layer.
@@
Input 9x6x1 9x6x1 -
Conv 1 9x6x1 9x6x4 LeakyReLU
IN 1 9x6x4 9x6x4 -
Conv 2 9x6x4 4x3x8 LeakyReLU
IN 2 4x3x8 4x3x8 -
Conv 3 4x3x8 2x1x16 LeakyReLU
IN 3 2x1x16 2x1x16 -
Conv 4 2x1x16 1x1x1 LeakyReLU
Ouput 1x1x1 1 ReLU
To ensure the creation of effective generators, the objectives for both generators and discriminators are established as follows:
@@
This loss function is inspired by generative adversarial networks (GANs) and encourages the generator to produce transformed data from one domain that is indistinguishable from real data in the other domain. It is computed by the discriminator network, which aims to classify the generated data as fake while the generator aims to fool the discriminator by generating realistic data.
$\begin{split}\mathcal{L}_{\text{adv}}(G,D_{Y},X,Y)=&\mathbb{E}_{y\sim p_{\text{data}}(y)}[\log D_{Y}(y)]+\\ &\mathbb{E}_{x\sim p_{\text{data}}(x)}[1-\log D_{Y}(G(x))]\end{split}$ (1)
$\begin{split}\mathcal{L}_{\text{adv}}(F,D_{X},Y,X)=&\mathbb{E}_{x\sim p_{\text{data}}(x)}[\log D_{X}(x)]+\\ &\mathbb{E}_{y\sim p_{\text{data}}(y)}[1-\log D_{X}(F(y))]\end{split}$ (2)
@@
The cycle consistency loss ensures that the mapping from the source domain to the target domain and back to the source domain is consistent. It measures the difference between the original input data and the data reconstructed after going through both generator mappings. Through the minimization of this loss, the preservation of crucial features during the transformation process is enforced.
$\mathcal{L}_{\text{cyc}}(G,F,X)=\mathbb{E}_{x\sim p_{\text{data}}(x)}[||x-F(G(x))||_{1}]$ (3)
$\mathcal{L}_{\text{cyc}}(F,G,Y)=\mathbb{E}_{y\sim p_{\text{data}}(y)}[||y-G(F(y))||_{1}]$ (4)
@@
$\mathcal{L}_{\text{id}}(G,X)=\mathbb{E}_{x\sim p_{\text{data}}(x)}[||G(x)-x||_{1}]$ (5)
$\mathcal{L}_{\text{id}}(F,Y)=\mathbb{E}_{y\sim p_{\text{data}}(y)}[||F(y)-y||_{1}]$ (6)
@@
The total loss is a combination of these losses, weighted by respective coefficients, and aims to optimize the generators and discriminators simultaneously to achieve high-quality time-series NDVI and EVI data transformation while maintaining consistency and identity preservation.
$\begin{split}\mathcal{L}_{\text{total}}=&\alpha(\mathcal{L}_{\text{adv}}(G,D_{Y},X,Y)+\mathcal{L}_{\text{adv}}(F,D_{X},Y,X))+\\ &\beta(\mathcal{L}_{\text{cyc}}(G,F,X)+\mathcal{L}_{\text{cyc}}(F,G,Y))+\\ &\sigma(\mathcal{L}_{\text{id}}(G,X)+\mathcal{L}_{\text{id}}(F,Y))\end{split}$ (7)
@@
In the evaluation of CycleGAN, it is common practice to assess the model’s performance on paired datasets, even though the model is trained without explicitly using paired data. However, in the case of CropGAN, there are no paired datasets available for evaluation due to the lack of labelled data in the target domain. This poses a challenge in evaluating the model’s performance using traditional paired dataset evaluation metrics.
To overcome this limitation, the total loss is employed as an evaluation metric. After the GropGAN network has reached a stable state for the training process after several training epochs, the model is selected associated with the smallest total loss. This particular model is then designated as our well-trained model.
This approach allows us to tackle the absence of paired datasets for evaluation in CropGAN and obtain an objective measure of the model’s performance.
IV-D Crop Mapper
TABLE IV: The Crop Mapper Structure. BN represents a batch normalization layer. Conv represents a convolutional layer. FC represents a fully connected layer.
Layer Input Shape Output Shape Activation
Input 9x6x1 9x6x1 -
Conv 1 9x6x1 9x6x2 ReLU
BN 1 9x6x2 9x6x2 -
Conv 2 9x6x2 4x3x2 ReLU
BN 2 4x3x2 4x3x2 -
Conv 3 4x3x2 2x1x4 ReLU
BN 3 2x1x4 2x1x4 -
Flatten 2x1x4 8 -
FC 1 8 4 ReLU
FC 2 4 1 ReLU
Output 1 1 Softmax
The crop mapper utilizes a Convolutional Neural Network (CNN) as the core component for crop classification. The CNN architecture consists of convolutional, pooling, and fully connected layers, designed to learn hierarchical representations of the input data and capture both low-level features and high-level semantic information. However, due to the small structure of the time-series remote sensing input data, pooling layers are omitted. The crop mapper structure is shown in Table IV.
A single CNN model is trained exclusively on the source domain data to optimize its performance for the specific characteristics and patterns of the source domain. As explained above, directly applying this trained model to the target domain can result in suboptimal performance due to domain shift, where the target domain data exhibits different distributions and variations compared to the source domain data. Therefore, before applying the trained crop mapper on a given target domain, the remote sensing data of the target domain should be first transformed using the domain mapper to resemble those of the source domain. Subsequently, to derive the crop mapping result, the transformed data is fed into the crop mapper.
By combining the domain mapper and the crop mapper, our methodology achieves robust and accurate crop mapping results even in the presence of domain shift. This approach leverages the strengths of both models to effectively tackle the challenges posed by different domains in crop mapping applications.

### v2

@@ III-B Reference Data
The CDL is used as the ground truth for the source domain. The CDL [7], a crop-specific land cover raster map dataset available for the entire conterminous U.S. land area (CONUS) at 30 m resolution provided by the USDA, regularly provides information on the annual temporal and spatial distribution of corn, as well as the area dedicated to its cultivation. However, for the target domain, official ground reference data for China is currently unavailable. Fortunately, [6] published maps of corn and soybean with a spatial resolution of 10 meters for Northeast China from 2017 to 2019. As a result, the 2019 crop maps from [6] are regarded as the ground truth reference for the study area in China.
@@ III-C Remote Sensing Data
As described, our primary objective is to locate the specific crop (corn) with cross-domain problems at an earlier growth stage. As a result, the time series remote sensing data should start after the planting period and conclude before the onset of the crop harvest period. Referring to Figure 3, it can be observed that the corn harvesting season typically starts in early September in the USA and begins in early October in China. Planting, on the other hand, starts at the beginning of April in the USA, and in mid-April in Jilin Province, China. As a result, the date of our annual collection of remote sensing data spans from May 1st to July 30th, encompassing a total of nine-time points with a ten-day observation window.
@@
Figure 3: The Crop Calendar for Corn. The red lines represent the average dates.
@@ IV Methodology
IV-A Problem Statement
Let $\mathbf{X}$ denote the time series remote sensing input data and $\mathbf{Y}$ denote the ground truth labels. Each sample $\mathbf{x}$ can be expressed as a temporal form $[\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{t}]$, where $\mathbf{x}_{i}$ represents input at time $i$. $\mathbf{x}_{i}$ can be further expanded as $[\mathbf{x}_{i1},\mathbf{x}_{i2},...,\mathbf{x}_{ib}]$, containing multi-spectral bands information from band $1$ to band $b$. Each ground truth label, denoted as $y$, is represented as a binary number, which corresponds to the categories of corn and other crops. Let $\mathbf{X}_{t}$ denote the time series input target data, $\mathbf{X}_{s}$ denote the time series input source data, and $\mathbf{Y}_{s}$ denote the GT for $\mathbf{X}_{s}$. Each sample $\mathbf{x}_{s}$ has a corresponding $\mathbf{y}_{s}$.
Our objective is to identify target crops to get target labels $\mathbf{Y}_{t}$ in the target domain during their early growth stages, utilizing labelled source domain data ($\mathbf{X}_{s}$, $\mathbf{Y}_{s}$) and unlabelled target domain data ($\mathbf{X}_{t}$). This approach addresses challenges related to the cross-domain issue and the large data distribution discrepancies issue.
IV-B System Overview
Figure 4: The Architecture of the CropSTGAN Framework.
The architecture of our CropSTGAN framework is depicted in Figure 4, showcasing both the training and mapping dataflows. This system includes three main components: the pre-processor, the CropSTGAN domain mapper, and the TempCNN crop mapper. During training, the pre-processor enhances MSI images quality by filling gaps through linear interpolation, ensuring complete MSI image series $\mathbf{X}_{s}$ and $\mathbf{X}_{t}$ for both source and target domains. The CropSTGAN domain mapper, trained with randomly sampled data from these MSI images, enables the transfer of time-series remote sensing data between domains: $\mathbf{X}_{t}\rightarrow\mathbf{X}_{s}$ and $\mathbf{X}_{s}\rightarrow\mathbf{X}_{t}$. The TempCNN crop mapper is then trained on labelled source domain data ($\mathbf{X}_{s}$, $\mathbf{Y}_{s}$) to identify the target crop in the source domain at an early stage.
For mapping, the process begins with the pre-processor supplying the target domain MSI image series $\mathbf{X}_{t}$. These images are then transformed by the CropSTGAN domain mapper from the target to the source domain as $\mathbf{X}_{s}^{{}^{\prime}}$. The trained TempCNN crop mapper uses these transformed images $\mathbf{X}_{s}^{{}^{\prime}}$ to locate the target crop $\mathbf{Y}_{t}$ in the target domain. This system design ensures accurate and reliable early-stage crop mapping in the target domain, particularly useful when labelled data is scarce or nonexistent, thereby overcoming the challenges of cross-domain mapping.
IV-C Pre-processor
@@
The pre-processor aims at providing complete time-series MSI data by filling gaps between MSI images due to cloud cover, atmospheric interference, or sensor limitations. The processing pipeline is outlined in Figure 5.
The procedure illustrated in the figure begins with the collection of MSI images covering the entire targeted study areas, sourced from the Sentinel-2 Dataset at consistent 10-day intervals. The primary objective is to detect the target crop at an early growth stage, addressing cross-domain challenges. Therefore, the selected time series of remote sensing data spans from May 1st to July 30th, starting post-planting and concluding pre-harvest, resulting in nine image sets. To refine the focus on agricultural lands, the method employs a cropland mask from the Dynamic World dataset during the crop’s growing season, adjusting it to a uniform 30-meter resolution. This step ensures the exclusion of non-agricultural areas, concentrating the analysis on crop areas within the nine groups of MSI images. For each set, cloud detection is performed using Sentinel Hub’s tool, filtering out cloud-covered areas. The process averages the values of cloud-free images to produce a composite MSI image with a clear 30-meter resolution. This approach yields nine high-quality, cloud-free time-series MSI images, each comprising six bands and maintaining a 30-meter spatial resolution.
IV-D CropSTGAN Domain Mapper
Figure 6: The CropSTGAN Domain Mapper Structure. (a) Cycle 1: The training process involves data flowing from the target domain to the source domain and then back to the target domain. The notation $\mathbf{X}_{s}^{{}^{\prime}}$ represents the transformed target domain data, while $\mathbf{X}_{t}^{{}^{\prime}}$ denotes the data transformed back to the target domain. (b) Cycle 2: The training process involves data flowing from the source domain to the target domain and then back to the source domain. The notation $\mathbf{X}_{t}^{{}^{\prime}}$ represents the transformed source domain data, while $\mathbf{X}_{s}^{{}^{\prime}}$ denotes the data transformed back to the source domain.
IV-D1 The Generators and Discriminators Structures
@@
Input 9$\times$6 9$\times$6 -
Encoder 1 9$\times$6 9$\times$1$\times$4 LeakyReLU
Encoder 2 9$\times$1$\times$4 7$\times$1$\times$8 LeakyReLU
Encoder 3 7$\times$1$\times$8 5$\times$1$\times$16 LeakyReLU
Encoder 4 5$\times$1$\times$16 3$\times$1$\times$32 LeakyReLU
Decoder 4 3$\times$1$\times$32 5$\times$1$\times$16 LeakyReLU
Decoder 3 5$\times$1$\times$16 7$\times$1$\times$8 LeakyReLU
Decoder 2 7$\times$1$\times$8 9$\times$1$\times$4 LeakyReLU
Decoder 1 9$\times$1$\times$4 9$\times$6$\times$4 LeakyReLU
Output 9$\times$6$\times$4 9$\times$6 ReLU
TABLE III: The Discriminator/Crop Mapper Structure. The last layer of the Discriminator uses ReLU as its activation function, while the Classifier employs Softmax.
@@
Input 9$\times$6 9$\times$6 -
Conv 1 9$\times$6 9$\times$1$\times$8 LeakyReLU
Conv 2 9$\times$1$\times$8 7$\times$1$\times$8 LeakyReLU
Conv 3 7$\times$1$\times$8 5$\times$1$\times$8 LeakyReLU
Conv 4 5$\times$1$\times$8 3$\times$1$\times$8 LeakyReLU
Flatten 3$\times$1$\times$8 24 -
FC1 24 4 ReLU
Output 4 1 ReLU or Softmax
The domain mapper is referred as CropSTGAN, which stands for Crop Mapping Spectral-temporal Generative Adversarial Network. CropSTGAN domain mapper serves as the key component in our framework for transforming time-series remote sensing data points between different domains. As shown in Figure 6, the domain mapper consists of two generator networks, $G$ and $F$, and two discriminator networks, $D_{X}$ and $D_{Y}$. In the training process, the generators and discriminators play distinct roles. The generators’ primary task is to transform the data from one domain to another: $G:\mathbf{X}_{t}\rightarrow\mathbf{X}_{s}^{{}^{\prime}}$ and $F:\mathbf{X}_{s}\rightarrow\mathbf{X}_{t}^{{}^{\prime}}$, while the discriminator’s main objective is to differentiate between real and transformed data. As the training progresses, both networks improve iteratively until an equilibrium is reached where the generator generates highly realistic data and the discriminator cannot reliably distinguish between real and fake samples. For example, when the discriminator $D_{s}$ can not distinguish whether the transformed data $\mathbf{X}_{s}^{{}^{\prime}}$ generated by the generator $G$ is from the source domain or not, it means that the generator $G$ has been well-trained.
The generator and discriminator structures are depicted in Table II and Table III, respectively. Each generator consists of four encoders and four decoders. Its first encoder layer, equipped with $4$ filters with a $3$ kernel size, captures temporal features across all 6 spectral bands. Subsequent encoders use the same kernel size but increase filter counts to 8, 16, and 32, respectively. The decoders, designed to mirror the encoders, reverse the encoding process, thereby reconstructing the original input features.
The discriminator shares a similar design with the generator’s encoder, with the distinction that its convolution layers uniformly utilize 8 filters. It then incorporates a flatten layer and a fully connected layer, with dimensions of 24 and 4, respectively. Its final layer is a fully connected layer with a single output, aimed at distinguishing between the transformed target domain data and the source domain data.
TABLE IV: The Generator Structure of CropTGAN. Each encoder layer consists of an instance normalization layer, an activation layer and a pooling layer. Each decoder layer consists of an instance normalization layer, an activation layer and an upsampling layer.
Layer Input Shape Output Shape Activation
Input 9$\times$6 9$\times$6 -
Encoder 1 9$\times$6 9$\times$6$\times$4 LeakyReLU
Encoder 2 9$\times$6$\times$4 7$\times$6$\times$8 LeakyReLU
Encoder 3 7$\times$6$\times$8 5$\times$6$\times$16 LeakyReLU
Encoder 4 5$\times$6$\times$16 3$\times$6$\times$32 LeakyReLU
Decoder 4 3$\times$6$\times$32 5$\times$6$\times$16 LeakyReLU
Decoder 3 5$\times$6$\times$16 7$\times$6$\times$8 LeakyReLU
Decoder 2 7$\times$6$\times$8 9$\times$6$\times$4 LeakyReLU
Decoder 1 9$\times$6$\times$4 9$\times$6$\times$4 LeakyReLU
Output 9$\times$6$\times$4 9$\times$6 ReLU
TABLE V: The Discriminator/Crop Mapper Structure of CropTGAN. The last layer of the Discriminator uses ReLU as its activation function, while the Classifier employs Softmax.
Layer Input Shape Output Shape Activation
Input 9$\times$6 9$\times$6 -
Conv 1 9$\times$6 9$\times$6$\times$8 LeakyReLU
Conv 2 9$\times$6$\times$8 7$\times$6$\times$8 LeakyReLU
Conv 3 7$\times$6$\times$8 5$\times$6$\times$8 LeakyReLU
Conv 4 5$\times$6$\times$8 3$\times$6$\times$8 LeakyReLU
Flatten 3$\times$6$\times$8 144 -
FC1 144 4 ReLU
Output 4 1 ReLU or Softmax
To assess the distinctive architecture of the CropSTGAN domain mapper, we conducted comparative experiments with a simpler yet analogous model named CropTGAN. This model features encoders with identical filter counts and sizes as those in CropSTGAN. The primary distinction between the generators and discriminators of the two models lies in the configuration of the first layers. Specifically, the CropTGAN generator and discriminator independently extract features from the temporal dimension on each band in their respective first encoder and convolutional layers. The details of the generators and discriminators are documented in Table IV and Table V.
IV-D2 Losses
As shown in Figure 6, there are two cycle training processes:
  • In the first cycle training process (Cycle 1), data from the target domain is first transformed to the source domain and then reverted back to the target domain: $G:\mathbf{X}_{t}\rightarrow\mathbf{X}_{s}^{{}^{\prime}}\rightarrow\mathbf{X}_{t}^{{}^{\prime}}$ and
  • In the second cycle training process (Cycle 2), data from the source domain is first transformed to the target domain and then reverted back to the source domain: $F:\mathbf{X}_{s}\rightarrow\mathbf{X}_{t}^{{}^{\prime}}\rightarrow\mathbf{X}_{s}^{{}^{\prime}}$
To ensure the creation of effective generators, the losses for both training processes are established as follows:
@@ IV-D2 Losses
This loss function is inspired by GANs and encourages the generator to produce transformed data from one domain that is indistinguishable from real data in the other domain. It is computed by the discriminator network, which aims to classify the generated data as fake while the generator aims to fool the discriminator by generating realistic data (source domain data).
$\begin{split}\mathcal{L}_{\text{adv1}}(G,D_{s},\mathbf{X}_{t},\mathbf{X}_{s})=&\mathbb{E}_{\mathbf{x}_{s}\sim p_{\text{data}}(\mathbf{x}_{s})}[\log D_{s}(\mathbf{x}_{s})]+\\ &\mathbb{E}_{\mathbf{x}_{t}\sim p_{\text{data}}(\mathbf{x}_{t})}[1-\log D_{s}(G(\mathbf{x}_{t}))]\end{split}$ (1)
$\begin{split}\mathcal{L}_{\text{adv2}}(F,D_{t},\mathbf{X}_{s},\mathbf{X}_{t})=&\mathbb{E}_{\mathbf{x}_{t}\sim p_{\text{data}}(\mathbf{x}_{t})}[\log D_{t}(\mathbf{x}_{t})]+\\ &\mathbb{E}_{\mathbf{x}_{s}\sim p_{\text{data}}(\mathbf{x}_{s})}[1-\log D_{t}(G(\mathbf{x}_{s}))]\end{split}$ (2)
@@
The cycle consistency loss ensures that the transformed data from one domain to another domain and back to the original domain is consistent with the original data. It measures the difference between the original input data and the data reconstructed after going through both generator mappings. Through the minimization of this loss, the preservation of crucial features during the transformation process is enforced.
$\mathcal{L}_{\text{cyc1}}(G,F,\mathbf{X}_{t})=\mathbb{E}_{\mathbf{x}_{t}\sim p_{\text{data}}(\mathbf{x}_{t})}[||\mathbf{x}_{t}-F(G(\mathbf{x}_{t}))||_{1}]$ (3)
$\mathcal{L}_{\text{cyc2}}(F,G,\mathbf{X}_{s})=\mathbb{E}_{\mathbf{x}_{s}\sim p_{\text{data}}(\mathbf{x}_{s})}[||\mathbf{x}_{s}-G(F(\mathbf{x}_{s}))||_{1}]$ (4)
@@
$\mathcal{L}_{\text{id1}}(G,\mathbf{X}_{t})=\mathbb{E}_{\mathbf{x}_{t}\sim p_{\text{data}}(\mathbf{x}_{t})}[||G(\mathbf{x}_{t})-\mathbf{x}_{t}||_{1}]$ (5)
$\mathcal{L}_{\text{id2}}(F,\mathbf{X}_{s})=\mathbb{E}_{\mathbf{x}_{s}\sim p_{\text{data}}(\mathbf{x}_{s})}[||F(\mathbf{x}_{s})-\mathbf{x}_{s}||_{1}]$ (6)
@@
The total loss is a combination of these losses, weighted by respective coefficients:
$\begin{split}\mathcal{L}_{\text{total}}=&\alpha(\mathcal{L}_{\text{adv1}}(G,D_{s},\mathbf{X}_{t},\mathbf{X}_{s})+\mathcal{L}_{\text{adv2}}(F,D_{t},\mathbf{X}_{s},\mathbf{X}_{t}))+\\ &\beta(\mathcal{L}_{\text{cyc1}}(G,F,\mathbf{X}_{t})+\mathcal{L}_{\text{cyc2}}(F,G,\mathbf{X}_{s}))+\\ &\sigma(\mathcal{L}_{\text{id1}}(G,\mathbf{X}_{t})+\mathcal{L}_{\text{id2}}(F,\mathbf{X}_{s}))\end{split}$ (7)
, where $\alpha,\beta,\sigma$ are the weight parameters. The total loss can be expressed as a minimax function:
$\min_{G,F}\max_{D_{t},D_{s}}\mathcal{L}_{\text{total}}(G,F,D_{t},D_{s},\mathbf{X}_{t},\mathbf{X}_{s})$ (8)
, where the generator seeks to minimize the loss while the discriminator aims to maximize it to achieve high-quality time-series MSI data transformation while maintaining consistency and identity preservation.
@@
The total loss is employed as an evaluation metric. After the GropSTGAN network has reached a stable state for the training process after several training epochs, the model is selected associated with the smallest total loss. This particular model is then designated as our well-trained model. This approach allows us to tackle the absence of paired datasets for evaluation in CropSTGAN and obtain an objective measure of the model’s performance.
IV-E TempCNN Crop Mapper
The TempCNN crop mapper denoted as $C$, structured similarly to the CropSTGAN domain mapper’s discriminators $D_{x}$ and $D_{y}$ as outlined in Table III, is trained solely with labelled source domain data ($\mathbf{X}_{s}$, $\mathbf{Y}_{s}$). The binary cross-entropy loss:
$\mathcal{L}_{\text{class}}(C)=-\sum_{s}\left(y_{s}\log(C(\mathbf{x}_{s}))+(1-y_{s})\log(1-C(\mathbf{x}_{s}))\right)$ (9)
, is used as the training loss for the TempCNN crop mapper. Its purpose is to locate the target crop within the source domain. However, applying this mapper directly in the target domain often leads to subpar results due to discrepancies in the variation of data distribution between the domains. To address this issue, the target domain’s MSI data must first be transformed to the source domain using the domain mapper’s generator. As illustrated in the right portion of Figure 4, this transformed data is then processed by the TempCNN crop mapper to generate mapping results:
$\mathbf{Y}_{t}=C(G(\mathbf{X}_{t}))$ (10)
This integrated approach ensures accurate and reliable early crop mapping in the target domain, even in the absence of the target domain ground truth data.